OpenStudy (anonymous):

Can someone please explain to me how to solve this. Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)). -9x - 2 -9x + 5 -9x + 25 -9x + 27

OpenStudy (anonymous):

replace the x in f(x) with -x + 3 f(g(x)) = 9(-x+3) - 2 = -9x +27 - 2 = -9x +25

jhonyy9 (jhonyy9):

nice @Orca !!!